OpenStudy (anonymous):

ok I have one more I think I am close to understanding this..5x+20=6x-5

OpenStudy (anonymous):

solving for x? x(5-6)=-(20+5) x=-(20+5)/(5+6)

OpenStudy (anonymous):

no I actually need the whole equation worked out

OpenStudy (anonymous):

what math is this?

OpenStudy (anonymous):

Well... Subtract 5x from both sides. and get 20=x-5. Then so a simple one-step equation.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

lol